# Navigating Rifts in Human-LLM Grounding: Study and Benchmark

### v1

@@ 5.2 Forecasting a User’s Grounding Acts
| Ambiguous | What causes tailbone pain? |
@@
Method
At turn , our goal is to forecast grounding act associated with user message . We achieve this by repurposing conditional training—concretely, we append each user message with a forecasting grounding token if the future turn contains a grounding act. We use our validated grounding acts labeler (§5.1) to obtain , using the high-level grounding categories as labels (e.g. , , and ).
We finetune Llama-3.1-8B on sequences of form where is the assistant response. At inference time, we provide and analyze the likelihoods (i.e. logits) of our tokens . We detail hyperparameters used in the training process in Appendix B. Beyond our finetuned forecaster, we also evaluate as a baseline few-shot prompted GPT-4o-mini (details in Appendix E.2).
Results
We find that our GPT-4o-mini few-shot baseline performs near-random, with Macro ROC AUC = 0.51. This result is not very surprising as forecasting is a challenging task: we must predict grounding acts without directly observing assistant responses . Our finetuned forecaster, however, performs significantly better (0.61). Full experimental results are in Appendix B. In the next section, we draw representative samples from the forecaster, constructing a benchmark where users are (un)likely to initiate a grounding process.
@@ 6 Rifts: A Grounding Benchmark
Dataset Details
Rifts consists of a final combined set of 1740 tasks (split counts in Appdx. Table 8). Each task in Rifts is a prompt drawn from an initial instruction in WildChat. Tasks are stratified based on how the user is predicted to continue the conversation: namely, are users predicted to grounding, a failure, or clarify an following an LLMs response (examples in Table 2). For these tasks, we would expect an LLM to take initiative—clarifying or following up appropriately. In addition, we include tasks where the user is expected to do none of the above, perhaps by switching the topic or ending the interaction altogether. Here, no grounding is required.
Curation Process
We construct Rifts by filtering prompts from WildChat, using the predicted grounding act of the forecaster from §5.2. Forecaster predictions inform us on whether a clarification or followup action will be required in the conversation. We first filtered correctly predicted tasks from forecasters trained444Why train on each split? To ensure fair evaluation, we create separate forecasters for train / val / test. This prevents leakage and enables researchers to develop interventions using the train forecaster before evaluating with the test forecaster. on each split (train / val / test). For each grounding category, we then extracted the top 150 tasks with the highest likelihood of generating advancing, addressing, or ambiguous forecasting tokens. In other words, we repurpose our forecaster to curate representative tasks across each grounding act, sorting by the logit associated with each forecasting token. In addition, we sample 150 tasks that have a high likelihood of generating no forecasting token, capturing tasks that do not require initiative. Finally, we apply basic quality controls (see Appendix C).
Evaluating LLMs
@@
Off-the-shelf models struggle
@@ 7 Discussion and Conclusion
Towards LLMs that Initiate Grounding
@@
Benchmarking Human-LLM Grounding
@@
We considered grounding and engagement with Bing Chat in the absence of access to existing system meta prompts. The system prompts can greatly shape the responses and provide specific guidance on the flow of dialogue. Rifts was also collected by filtering WildChat using our forecaster; therefore, Rifts will only reflect tasks seen in WildChat. In addition, tasks in Rifts also depend on the LLMs used to serve WildChat (e.g. OpenAI LLMs). More specifically, our forecasters implicitly learn what tasks fail for the GPT models deployed in WildChat. Regardless, we observe that our final Rifts tasks are challenging for all evaluated models. Finally, our annotator relies on GPT-4o-mini to label logs with grounding acts. While we did show that the annotator generally agrees with human judgment on a subset of the data (§5.1), the annotator is not perfect.
@@
Enhancing an LLM’s ability to generate grounding acts (by initiating clarification and follow-up actions) raises potential concerns around privacy, as these actions may lead users to disclose sensitive information unintentionally. Balancing the need for grounding with the careful collection of only relevant information remains a significant challenge and an area for future research. Moreover, while effective grounding can improve interaction quality, it can also be misused in harmful contexts. Although our work focuses on improving grounding for constructive purposes, such as assisting users, these techniques could be exploited for harmful ends (e.g., manipulation or coercion in sensitive areas like political targeting).
Finally, our description of human-LLM grounding does not imply that LLMs possess genuine understanding. Like prior work, we use grounding as a metaphor to describe interaction processes that help align human expectations with LLM-generated responses Paek and Horvitz (2003); Brennan (2014); Shaikh et al. (2024). While human conversation involves active mutual comprehension, the same cannot be said of LLMs. The use of grounding terminology in this work is intended as a conceptual tool to analyze how LLMs facilitate or hinder effective communication, not as an anthropomorphic assertion that they share human-like cognitive capacities.
@@ Acknowledgments
We thank members of the HAX team, The SALT Lab, Will Epperson, Michael Li, Jan-Philipp Fränken, Michael Ryan, Yanzhe Zhang, Qinan Yu, Shardul Sapkota and Ryan Li.
@@ A.2 Descriptive Analysis
| Ambiguous | bitcoin (4.66), chatbot (4.24), cryptocurrency (4.01), unauthorized (2.39), beginners (2.19), friends (2.02) |
@@ B.3 Reweight Control Tokens At Train Time
We want to encourage our model to learn our added forecasting tokens alongside the language modeling objective. The standard MLE objective optimizes a model’s parameters with respect to a sequence : . However, a subsequence consists of forecasting tokens, which we want to emphasize—especially since these tokens did not undergo pretraining. At train time, we upweight these tokens by . Our final loss is below:
@@ Appendix D Rifts
| Model | Addressing (%) | Advancing (%) | No Grounding (%) | Ambiguous (%) |

### v2

@@ 5.2 Forecasting Users’ Grounding Acts
| Disambiguating | What causes tailbone pain? |
@@
Method.
Following a user message , our goal is to forecast grounding act associated with the next user message . We achieve this by repurposing conditional training. Concretely, we append each user message with a forecasting grounding token if the future turn contains a grounding act. Consider the hypothetical training example below:
In the example above, we append a new forecasting token after the user’s initial message , since the user’s following turn is a repair. We use our validated grounding acts labeler (§5.1) to obtain all , using the high-level grounding categories as labels (e.g. , , and ). We finetune Llama-3.1-8B on sequences of form where is the assistant response.
At inference time, we can provide any user message and analyze the predicted likelihoods (i.e. logits) of our grounding tokens that are generated right after. For example:
This enables us to predict—from just a user query alone—the user’s predicted grounding outcome independent of the model’s response! This is an especially challenging learning problem: we effectively marginalize over all possible assistant responses. Successfully learning this model enables early intervention. All hyperparameters used in the training process are outlined in Appendix B. Beyond our finetuned forecaster, we also evaluate as a baseline few-shot prompted GPT-4o-mini (details in Appendix E.2).
Results.
We find that our GPT-4o-mini few-shot baseline performs near-random, with Macro ROC AUC = 0.51. This result is not very surprising as forecasting is a challenging task: we must predict grounding acts without directly observing assistant responses . Our intuition is that it’s much easier to look at an entire conversation and label the conversation for grounding acts post-hoc than it is to forecast a likely grounding act without being able to “see” future turns. Our finetuned forecaster, however, performs significantly better (0.61). Full experimental results are in Appendix B. In the next section, we draw representative samples from the forecaster, constructing a benchmark where users are (un)likely to initiate a grounding process.
@@ 6 Rifts: A Grounding Benchmark
Dataset Details.
Rifts consists of a final combined set of 1740 tasks (split counts in Appdx. Table 8). Each task in Rifts is a prompt drawn from an initial instruction in WildChat. Tasks are stratified based on how the user is predicted to continue the conversation: namely, are users predicted to grounding, a failure, or following an LLMs response (examples in Table 2). For these tasks, we would expect an LLM to take initiative—clarifying or following up appropriately. In addition, we include tasks where the user is expected to do none of the above, perhaps by switching the topic or ending the interaction altogether. Here, no grounding is required.
Curation Process.
We construct Rifts by filtering prompts from WildChat, using the predicted grounding act of the forecaster from §5.2. Forecaster predictions inform us on whether a clarification or followup action will be required in the conversation. In building Rifts, we implicitly hypothesize that for some prompts , regardless of what the LLM replies with (), the user is in grounding trouble. In other words, if a user gives a query to a model that’s so severely underspecified (e.g. = “write me a resume”), it does not matter what any LLM responds with. The user must go back and forth to build common ground, since they never gave enough information in the initial prompt. Rifts identifies this class of prompts, using the forecaster as a proxy.
We first filtered correctly predicted tasks from forecasters trained444Why train on each split? To ensure fair evaluation, we create separate forecasters for train / val / test. This prevents leakage and enables researchers to develop interventions using the train forecaster before evaluating with the test forecaster. on each split (train / val / test). For each grounding category, we then extracted the top 150 tasks with the highest likelihood of generating advancing, addressing, or ambiguous forecasting tokens. In other words, we repurpose our forecaster to curate representative tasks across each grounding act, sorting by the logit associated with each forecasting token. In addition, we sample 150 tasks that have a high likelihood of generating no forecasting token, capturing tasks that do not require initiative. Finally, we apply basic quality controls (see Appendix C).
Evaluating LLMs.
@@
Off-the-shelf models struggle.
@@ 7 Discussion and Conclusion
Should we expect grounding behavior from LLMs?
Perhaps we should not be surprised that LLMs are unable to initiate grounding, defaulting instead to instruction-following. Models that are not trained to follow instructions are already biased towards instruction following behavior, likely because of the large presence of instruction following articles in pre-training mixes Hewitt et al. (2024). In addition, limitations in theory of mind and other metacognitive challenges may restrict the ability of models to engage in grounding interactions Sap et al. (2022); Ullman (2023). Training methods must counteract these limitations and biases. Still, we see promise in future methods that elicit grounding capabilities from LLMs; and Rifts can serve as a resource to test these methods.
Towards LLMs that initiate grounding.
@@
Benchmarking human-LLM grounding.
@@
We considered grounding and engagement with Bing Chat in the absence of access to existing system meta prompts. System prompts can greatly shape the responses and provide specific guidance on the flow of dialogue. Rifts was also collected by filtering WildChat using our forecaster; therefore, Rifts will only reflect tasks seen in WildChat. In addition, tasks in Rifts also depend on the LLMs used to serve WildChat (e.g. OpenAI LLMs). More specifically, our forecasters implicitly learn what tasks fail for the GPT models deployed in WildChat. Regardless, we observe that our final Rifts tasks are challenging for all evaluated models. Finally, our annotator relies on GPT-4o-mini to label logs with grounding acts. While we did show that the annotator generally agrees with human judgment on a subset of the data (§5.1), the annotator is not perfect.
@@
Enhancing an LLM’s ability to generate grounding acts (by initiating clarification and follow-up actions) raises potential concerns around privacy, as these actions may lead users to disclose sensitive information unintentionally. Balancing the need for grounding with the careful collection of only relevant information remains a significant challenge and an area for future research. Moreover, while effective grounding can improve interaction quality, it can also be misused in harmful contexts. Although our work focuses on improving grounding for constructive purposes, such as assisting users, these techniques could be exploited for harmful ends (e.g., manipulation, persuasion, or coercion in sensitive areas like political targeting).
Finally, our description of human-LLM grounding does not imply that LLMs possess genuine understanding. Like prior work, we use grounding acts to describe interaction processes that help align human expectations with LLM-generated responses Paek and Horvitz (2003); Brennan (2014); Shaikh et al. (2024). While human conversation involves active mutual comprehension, the same cannot be said of LLMs. The use of grounding terminology in this work is intended as a conceptual tool to analyze how LLMs facilitate or hinder effective communication, not as an anthropomorphic assertion that they share human-like cognitive capacities.
@@ Acknowledgments
We appreciate the feedback from members of the Microsoft HAX team, the Stanford SALT Lab, Will Epperson, Michael Li, Jan-Philipp Fränken, Michael Ryan, Yanzhe Zhang, Qinan Yu, Shardul Sapkota and Ryan Li.
@@ A.2 Descriptive Analysis
| Disambiguating | bitcoin (4.66), chatbot (4.24), cryptocurrency (4.01), unauthorized (2.39), beginners (2.19), friends (2.02) |
@@ B.3 Reweight Control Tokens At Train Time
We seek to encourage our model to learn our added forecasting tokens alongside the language modeling objective. The standard MLE objective optimizes a model’s parameters with respect to a sequence : . However, a subsequence consists of forecasting tokens, which we want to emphasize—especially since these tokens did not undergo pretraining. At training time, we upweight these tokens by . Our final loss is below:
@@ Appendix D Rifts
| Model | Addressing (%) | Advancing (%) | No Grounding (%) | Disambiguating (%) |